princeevee:

Adam has 10 coins in his pocket. He has 2 nickels, 3 quarters, 2 dimes and 3 pennies. Adam chooses a coin at random. What is the probability that he selects a nickel or a dime from his pocket?

princeevee:

@Hero

Hero:

The question type is the same as the previous question which means the process to find the answer is the same except the numbers have changed.

princeevee:

4/10 is the fraction, a.k.a its 2 x 2 /10

princeevee:

2/5?

Hero:

Show all of your work please. How did you come up with 4/10 to begin with?

princeevee:

well, 2 dimes and nickles out of all the coins, which is 10, its really simple that i dont need to actually explain

Hero:

The answer is correct. You executed nearly flawlessly but you posted something erroneous. You must add the fractions 2/10 + 2/10 = (2 + 2)/10 = 4/10 = 2/5